$48000 cost, $8000 down payment and so $40000 of loan and 6 years, 5%. How much money you need to pay at the end?



And there is a example

2
media%2F1c1%2F1c111b09-15f6-4ea7-bd37-e3
Answer #1

Loan amount = 40,000,   i = 5%,   n = 6 years

Future Value = P(F/P, i, n)

                    = 40,000(F/P, 5%, 6)

                    = 40,000(1.340)

                   = $53,600

Thus, you need to pay $53,600 at the end.